OpenStudy (anonymous):

Find a and b correct to the nearest hundredth using the Pythagorean Theorem. 1) a = 3b, c = 20 2) a = 2/3b, c = 52

OpenStudy (anonymous):

(3b)^2 +b^2 =20^2 9b^2 +b^2 =400 10b^2=400 b^2=40 b=sqrt(40)

OpenStudy (anonymous):

since a=3b a=3sqrt(40)